OpenStudy (boobear7411):

Factor the following polynomials. (1 point each) 9. 3x2 – 12 10. 5x3y + 35x2y + 45xy

OpenStudy (anonymous):

I'm assuming that 1 point each is supposed to be your points, so I don't want to steal them from you. When factoring a polynomial you want to get two factors like (x+a)(x+b) where a, b are integers. It doesn't always work out well but these problems are designed so that they do factor nicely. So let's look at the first one: 3x^2 - 12 A polynomial like this (second degree, meaning highest exponent is 2) should look like: Ax^2 + Bx + C where A, B, C are all integers and may or may not be negative. In this example, we can divide the expression by 3 because 12/3 is easy and 3/3 is even easier. This gives: 3(x^2 - 4). Here's a neat little trick: (x + a)(x - a) = x^2 + a^2 Applying that here: 3(x^2 + 4) = 3(x^2 + 2^2) = 3(x + 2)(x - 2). And we're done! Now you do the second one.

OpenStudy (anonymous):

Oops I meant (x + a)(x - a) = x^2 - a^2

OpenStudy (boobear7411):

Thank You sooooo much